OpenStudy (anonymous):

One number is four less than three times another. If their sum is increased by five, their result is twenty five. Find the numbers

OpenStudy (anonymous):

Let one number=x and the other number=y Condition 1:One number is four less than three times another x = (y-4) (1st Equation) Condition 2:If their sum is increased by five, their result is twenty five (x + y) + 5 = 25 (2nd Equation) Solving both equations we get, x = 8 and, y = 12.